geerky42 (geerky42):

A parabola has roots at 0.4 and -3. The equation is f(x) = 6-ax-bx². Find the value of b.

OpenStudy (anonymous):

product of roots =constant/coeff. of x^2 so, 0.4*-3=-1.2=6/-b b=6/1.2=5 :)

geerky42 (geerky42):

Wut? How come the product of roots = constant/coefficient of x² ?

geerky42 (geerky42):

@Aperogalics

OpenStudy (anonymous):

@geerky42 can u answer my simple answers?????????

geerky42 (geerky42):

What do you mean?

OpenStudy (anonymous):

i mean to say that give me answers of some simple questions related to this and u will get to know the answer.............r u ready?????????

geerky42 (geerky42):

Sure.

OpenStudy (anonymous):

ok 0.4 and -3 are roots so (x-0.4) and (x-(-3)) are factors.....right

geerky42 (geerky42):

Yeah, (x-0.4)(x+3) = f(x) f(x) is also equal to 6-ax-bx², right?

OpenStudy (anonymous):

yup:) now see this